OpenStudy (anonymous):

Could someone go step by step explaining how to simplify the expression and then add or subtract so that I may take notes and possibly understand how to do the problem? \[\sqrt{3x^3}+x\sqrt{75x}-\sqrt{27^x3}]\]\[\sqrt{3x^3} + [x\sqrt{75x}] -[\sqrt{27^x3}]

OpenStudy (anonymous):

The last term does not match the others, could there be a typo in your question?

OpenStudy (anonymous):

there is not supposed to be a square bracket at the end

OpenStudy (anonymous):

oh and the X is just an X not up high like that

OpenStudy (anonymous):

the 3 is supposed to be the exponent not the X

OpenStudy (anonymous):

I'm sorry I flubbed up trying to type it in

OpenStudy (anonymous):

\[ \sqrt{3x^3}+x\sqrt{75x}+\sqrt{27x^3}\] \[ \sqrt{x^23x}+x\sqrt{25(3)x}+\sqrt{9x^23x}\] \[ x\sqrt{3x}+5x\sqrt{3x}+3x\sqrt{3x}\] \[9x\sqrt{3x}\]

OpenStudy (anonymous):

mark on the first part in simplifying you break up the ^3 to simplify? I am sorry in advance but this is really hard for me

OpenStudy (anonymous):

We have square roots so we want to pull out squares.

OpenStudy (anonymous):

In the second term we write 75 as (5^2)(3) and the 5^2 can come out from under the radical as 5.